Calculation of cost of capital

The company’s CFO has informed the project team that the company plans to raise new capital to fund the investment in the project. The company’s current capital structure
consists of the following:

Debt : 700,000 7.4% coupon secured notes with five years to maturity. These notes are currently priced at 95% of face value of $100, with half yearly coupon payments.

Equity: 83,300,000 ordinary shares outstanding with a par value of $1 and currently selling for $2.00 per share.

Market: the company’s ordinary shares currently have a beta of 1.1, the market risk premium is 7%, and the risk-free rate is 3.5%.

The CFO has indicated that the company would use its investment bank Development Finance to raise all the capital required for the project, should it be approved, consistent
with its current capital structure. Moreover, the company would not be able to allocate any retained earnings to the project as these funds are already committed to other projects.
Subsequent discussions with Development Finance indicated that the fees associated with raising new capital would be 8% on an issue of ordinary shares, and 4% on new debt issues. Development Finance has specified that these fees cover all direct and indirect issue costs.

Tax rate is 30%

Calculate cost of capital for project?
Answer #1

Answer:

Market value of Debt = 700,000 * $95 = $ 66.5 mn

Market value of Equity = 83,300,000 shares * $2 per share = $ 166.6 mn

Therefore,

Debt (as % of Total Capital) = 66.5 / (66.5 + 166.6) = 28.53%, and

Equity (as % of Total Capital) = 166.6 / (66.5 + 166.6) = 71.47%

Now, bond that pays 2 coupon(s) of 7.40% per year, that has a market value of $95, and that matures in 5 years will have a yield to maturity of 4.33%. This implies a pre-tax annualized Cost of Debt of 8.85%

Cost of Equity = Risk-free rate + beta * market premium = 3.50% + 1.1 * 7% = 11.20%

Effective cost of Debt (post-tax) = 8.85% * (1.04) * (1-30%) = 6.44%

Effective cost of Equity = 11.20% * (1.08) = 12.10%

Cost of Capital = Proportion of Debt * Effective Cost of Debt + Proportion of Equity * Effective Cost of Equity = 28.53% of 6.44% + 71.47% of 12.10% = 10.48%.